1÷(secA-tan A)=secA+tanA how to prove,....,.....................

Grade:11

1 Answers

Arun
6 years ago
we know by identity
sec^2 A –tan^2 A= 1
(secA+tanA) (secA –tanA) = 1    by formula of a2 –b2= (a+b)(a –b)
hence secA+ tanA= 1/(secA –tanA)
hence proved